# Scale-DiT: Ultra-High-Resolution Image Generation with Hierarchical Local Attention

### v1

@@ Abstract
Ultra-high-resolution text-to-image generation demands both fine-grained texture synthesis and globally coherent structure, yet current diffusion models remain constrained to sub-$1K\times 1K$ resolutions due to the prohibitive quadratic complexity of attention and the scarcity of native $4K$ training data. We present Scale-DiT, a new diffusion framework that introduces hierarchical local attention with low-resolution global guidance, enabling efficient, scalable, and semantically coherent image synthesis at ultra-high resolutions. Specifically, high-resolution latents are divided into fixed-size local windows to reduce attention complexity from quadratic to near-linear, while a low-resolution latent equipped with scaled positional anchors injects global semantics. A lightweight LoRA adaptation bridges global and local pathways during denoising, ensuring consistency across structure and detail. To maximize inference efficiency, we repermute token sequence in Hilbert curve order and implement a fused-kernel for skipping masked operations, resulting in a GPU-friendly design. Extensive experiments demonstrate that Scale-DiT achieves more than $2\times$ faster inference and lower memory usage compared to dense attention baselines, while reliably scaling to $4K\times 4K$ resolution without requiring additional high-resolution training data. On both quantitative benchmarks (FID, IS, CLIP Score) and qualitative comparisons, Scale-DiT delivers superior global coherence and sharper local detail, matching or outperforming state-of-the-art methods that rely on native 4K training. Taken together, these results highlight hierarchical local attention with guided low-resolution anchors as a promising and effective approach for advancing ultra-high-resolution image generation.
Figure 1: Ultra-high-resolution images generated by Scale-DiT at $4K\times 4K$, $2K\times 3K$, and $2K\times 2K$. Zoomed-in regions highlight the fine-grained details preserved at these scales.
@@ 1 Introduction
Ultra-high-resolution image generation is becoming increasingly crucial for both creative and practical applications. From digital art and advertising to scientific visualization and virtual environments, users demand outputs with fine-grained textures, faithful structures, and seamless details at $4K$ resolution or higher. However, despite the rapid progress of text-to-image (T2I) diffusion models Ding et al. (2021); Rombach et al. (2022); Chen et al. (2023); Podell et al. (2023); Peebles & Xie (2023); Xue et al. (2023); Li et al. (2024b); Chen et al. (2024b); Li et al. (2024a); Black Forest Labs (2025); Cai et al. (2025); Xie et al. (2024; 2025), most state-of-the-art systems remain confined to resolutions below $1K\times 1K$. This limitation fundamentally restricts their usability in scenarios where ultra-high fidelity is paramount.
@@
The root of this challenge lies in two interdependent bottlenecks. First, scaling model capacity and datasets to higher resolutions requires tremendous resources, as high-quality 2K2K8K8K training data are scarce and expensive to curate. Second, attention-based diffusion architectures suffer from quadratic growth in token complexity with respect to image resolution, making naive scaling computationally prohibitive. For example, generating a $4K$ image would require tens of thousands of tokens, leading to impractical memory and runtime costs. As a result, existing approaches either resort to costly retraining on synthetic high-resolution datasets Hoogeboom et al. (2023); Liu et al. (2024a); Ren et al. (2024); Teng et al. (2023); Zheng et al. (2024); Zhang et al. (2025b), or adopt training-free upscaling methods Guo et al. (2024); Qiu et al. (2024); Du et al. (2024b); Liu et al. (2024b); Wu et al. (2025); Shi et al. (2025); Kim et al. (2025); Huang et al. (2024); Bu et al. (2025) that often compromise efficiency and stability. Consequently, the field still lacks a solution that delivers both ultra-high-resolution fidelity and computational scalability.
@@
In this work, we introduce Scale-DiT, a novel framework that rethinks ultra-high-resolution synthesis through the lens of efficiency and scalability. Inspired by the way artists construct large-scale murals, we decompose the generation task hierarchically: local windows capture fine textures through lightweight attention, while a low-resolution guidance image preserves global semantic structure. Technically, Scale-DiT introduces three innovations. First, a hierarchical local attention mechanism restricts computation to fixed-size windows, reducing quadratic cost to near-linear scaling while yielding over $10GB$ memory savings and more than $2\times$ faster inference compared to dense attention. Second, a global guidance pathway employs a low-resolution latent with scaled RoPE positional anchors to maintain long-range dependencies and ensure semantic coherence across windows. Third, a parameter-efficient joint denoising framework integrates global and local pathways via LoRA-adapted projections trained only on $256\times 256$$1K\times 1K$ data, enabling direct scaling to $4K$ or higher without requiring any native high-resolution training. These design choices translate into two key outcomes: (1) scalable fidelity, achieved by hierarchically fusing global composition with local detail to extend beyond pretrained resolutions, and (2) computational efficiency, enabled by reducing computation, memory, and runtime while reusing pretrained weights with lightweight adaptation.
@@
Extensive experiments validate these contributions: Scale-DiT reliably scales to $4K\times 4K$ synthesis with commodity training resolutions, achieving sharper textures, richer details, and stronger global coherence than existing approaches. Quantitatively, it surpasses dense attention baselines with over $2\times$ faster inference and lower memory usage, and it matches or outperforms state-of-the-art methods trained with native 4K data across FID, IS, and CLIP Score. Together, these results highlight Scale-DiT as a practical and effective solution for advancing ultra-high-resolution text-to-image generation.
@@
  • We demonstrate substantial efficiency improvements ($2\times$ speedup, less memory usage) while achieving state-of-the-art quality at $4K$, offering a practical and generalizable solution for real-world ultra-high-resolution generation.
@@ 2 Related Work
Text-to-Image Generation. Recent years have witnessed rapid progress in text-to-image (T2I) generation Ding et al. (2021); Rombach et al. (2022); Chen et al. (2023); Podell et al. (2023); Peebles & Xie (2023); Xue et al. (2023); Li et al. (2024b); Chen et al. (2024b); Li et al. (2024a); Black Forest Labs (2025); Cai et al. (2025); Xie et al. (2024; 2025); Gao et al. (2025), where diffusion models now achieve near-photorealistic synthesis at resolutions up to $1K\times 1K$. The dominant paradigm relies on cross-attention mechanisms and DiT architecture Peebles & Xie (2023), as seen in the PixArt series Chen et al. (2023; 2024b; 2024a), or multi-modal diffusion transformers (MMDiT) such as FLUX.1.0-dev Black Forest Labs (2025). These architectures have established strong foundations for controllable and high-quality T2I synthesis. However, despite their success, existing methods struggle to scale beyond pretrained resolutions due to quadratic attention costs and the lack of large-scale high-resolution data. Our work builds directly on this line of research, but departs from the prevailing direction by focusing on resolution scalability and computational efficiency rather than retraining larger models.
@@
High-resolution image synthesis. Real-world applications increasingly demand resolutions of $4K$ or higher, sparking intense research interest in pushing beyond the $1K\times 1K$ barrier. Several works, including PixArt-$\Sigma$ and SANA 1.5 Xie et al. (2025), have achieved near-$4K$ synthesis through extensive high-resolution pretraining, while others Hoogeboom et al. (2023); Liu et al. (2024a); Ren et al. (2024); Xie et al. (2023); Teng et al. (2023); Zheng et al. (2024); Zhang et al. (2025b); Yu et al. (2025) pursue fine-tuning or training-from-scratch approaches using curated datasets. For example, Zhang et al. (2025b) used wavelet supervision to enhance detail clarity, and Yu et al. (2025) proposed lightweight fine-tuning for adapting to higher resolutions. Although effective, these methods remain constrained by the scarcity of high-quality high-resolution data and substantial GPU requirements. More recent training-free strategies Guo et al. (2024); Qiu et al. (2024); Du et al. (2024b); Liu et al. (2024b); Wu et al. (2025); Shi et al. (2025); Kim et al. (2025); Huang et al. (2024); Bu et al. (2025) avoid data collection by leveraging pretrained models directly. For instance, I-Max Du et al. (2024b) aligns high-resolution flow with low-resolution manifolds, and HiFlow Bu et al. (2025) introduces low-resolution initialization to guide denoising. While promising, these approaches often inherit significant runtime and memory overhead, limiting accessibility for broader use. In contrast, our method unifies the strengths of data-free guidance and efficiency: by combining low-resolution references with local window attention, we achieve ultra-high-resolution synthesis without requiring new datasets or excessive hardware.
[…]
@@
Figure 2: Schematic of Scale-DiT’s attention block modifications. The left column illustrates that high-resolution image latents $X$ are partitioned into local windows (red) in Hilbert-Curve order that attend to each other within their window. Simultaneously, a low-resolution guidance latent $X_{lr}$ (yellow) provides global context to each window via position scaling. The right column shows the joint-denoising process and the attention kernel, that $X,X_{lr}$ are processed together with LoRA applied on the $X_{lr}$ part. The attention mask (upper right) enforces local and guidance-specific interactions and tile skipping in attention calculation to enable efficient and coherent ultra-high-resolution generation. No additional high-resolution training data are needed.
@@ 3 Method
Figure 2 illustrates the overall framework of Scale-DiT, highlighting the key modifications introduced in the attention blocks to support ultra-high-resolution generation. High-resolution image latents $X$ are partitioned into local windows (red grids), where tokens attend only within their window to capture fine-grained detail efficiently. A low-resolution guidance latent $X_{lr}$ (yellow), enhanced with scaled positional anchors, injects global semantics and preserves long-range consistency across windows. The attention mask is designed to enforce both local and guidance-specific interactions while also enabling tile-skipping to avoid unnecessary computation. Together, these mechanisms allow Scale-DiT to generate coherent and detail-rich $4K$ images with near-linear complexity and without requiring any native high-resolution training data. This schematic provides the foundation for the following technical components.
The Multi-Modal Diffusion Transformer (MMDiT) Preliminaries. The MMDiT employed in state-of-the-art models Black Forest Labs (2025); Cai et al. (2025); Gao et al. (2025), represents the current benchmark architecture for text-to-image generation. MMDiT processes two distinct token modalities: a noisy image token sequence $X\in\mathbb{R}^{N\times d}$ and a text token sequence $C_{T}\in\mathbb{R}^{M\times d}$. MMDiT processes image and text tokens through a unified Multi-Modal Attention (MMA) mechanism. Specifically, image and text tokens are concatenated as $C_{T};X]$ and processed via self-attention.
Within the MMA framework, spatial information is encoded using Rotary Position Embedding (RoPE) Su et al. (2024), which applies rotational transformations to capture relative positional relationships. This is mathematically represented as: $X_{i,j}=X_{i,j}\cdot\text{Rot}(i,j)$, where $\text{Rot}(i,j)$ denotes the rotation matrix corresponding to position $(i,j)$ in the 2D spatial grid. The MMA mechanism is formally defined as:
$\text{MMA}([C_{T};X])=\text{Softmax}\left(\frac{\left([Q_{T}({C_{T}}),Q(X)]\right)([K_{T}(C_{T}),K(X)]^{T})}{\sqrt{d}}\right)[V_{T}(C_{T}),V(X)]$
@@
Efficient Local Window Attention. Transformer-based generative models rely on self-attention, whose computational and memory costs scale quadratically with the number of tokens. For an image of size $H\times W$, the number of spatial tokens is $N=H\cdot W$, leading to $O(N^{2})$ complexity. Scaling from $1024\times 1024$ ($4096$ tokens [^1] A VAE downsampling factor of 16 yields $\frac{1024}{16}\times\frac{1024}{16}=4096$ tokens, similar scaling for other resolutions.) to $4096\times 4096$ ($65536$ tokens) increases the cost by a factor of $256$, which is prohibitive even with optimized kernels such as FlashAttention Dao et al. (2022); Dao (2023); Shah et al. (2024).
@@
To overcome this limitation, we partition the high-resolution latent $X\in\mathbb{R}^{H\times W}$ into non-overlapping windows $x_{i}\in\mathbb{R}^{l\times l}$}, with window size $l$ bounded by the pretrained resolution (e.g., 1024). Self-attention is computed independently within each window, reducing the complexity from $O(N^{2})$ to $O\left(\lceil\tfrac{H}{l}\rceil\cdot\lceil\tfrac{W}{l}\rceil\cdot(l^{2})^{2}\right)\approx O(N\cdot l^{2})$. In practice, we set $l=16$ (corresponding to $256\times 256$ windows) to strike a balance between accuracy and efficiency: larger windows bring little additional benefit in quality, while smaller windows underutilize GPU kernels, which are optimized for tile sizes around $128\times 128$ [^2] With 16× VAE downsampling, 256×256 windows yield 16×16 tokens, creating attention matrices that efficiently utilize GPU kernels. Smaller window size produces tokens less than the granularity of 128.. An ablation study in Section 5 further analyzes this trade-off.
With this design, scaling to $4K$ resolution ($N=65536$ tokens) reduces the attention complexity from $O(65536^{2})$ to $O(65536\cdot 16^{2})$, yielding a $256\times$ reduction. To maintain boundary consistency, we also allow limited cross-window attention along adjacent regions. Crucially, by fixing the window size regardless of the overall resolution, the local window cost remains constant as $H$ and $W$ grow. This makes the runtime scale nearly linearly with image resolution, enabling efficient and memory-friendly ultra-high-resolution generation while preserving fine-grained detail within each window.
@@
Maintaining Global Semantics via Low-Resolution Guidance. Partitioning a high-resolution image into local windows risks creating discontinuities and losing global semantic coherence. The key insight is that global long-range dependencies primarily influence the overall structure and layout of the generated image, but have minimal effect on fine local details. We generate a low-resolution guidance image $X_{lr}\in\mathbb{R}^{h\times w}$ to provide global context. The corresponding position indices are denoted as (m,n){(m,n)}, where $m\in\{0,1,\dots,h-1\}$ and $n\in\{0,1,\dots,w-1\}$. We define the scaling ratio as $\rho=\frac{H}{h}$ (empirically set to 4 for an optimal balance of performance and efficiency). We then scale the low-resolution image’s position indices by this ratio, mapping them to $(\tilde{m},\tilde{n})=(\rho\cdot m,\rho\cdot n)$. This effectively projects $X_{lr}$ to the same spatial scale as the high-resolution image $X$. Each token $X_{lr}[\tilde{m},\tilde{n}$] acts as an anchor point, providing contextual information. Each high-resolution window attends to its local tokens and the corresponding scaled region in $X_{lr}$, while $X_{lr}$ tokens attend globally among themselves and to text tokens. This ensures semantic consistency across distant regions. The framework naturally supports recursive generation: the high-resolution output can serve as low-resolution guidance for an even higher resolution, enabling arbitrarily large-scale synthesis with stable memory and computation.
@@
Parameter-Efficient Joint Denoising. To integrate the low-resolution guidance $X_{lr}$, we concatenate it with the standard text–image sequence to form $C_{T},X,X_{lr}]$. Since $X_{lr}$ is still an image modality, we can reuse the pretrained VAE and DiT blocks for its processing. However, scaling the positional indices of $X_{lr}$ alters the frequency characteristics of the RoPE embeddings. To adapt to this change, we fine-tune only the query, key, and value projections ($Q,K,V$) that process $X_{lr}$ using LoRA Hu et al. (2022), yielding adapted matrices $\tilde{Q},\tilde{K},\tilde{V}$. Importantly, the original $Q,K,V$ are still used for the high-resolution windows $X$, ensuring that pretrained generative capabilities for local content are preserved. The unified attention mechanism, simplified by omitting text tokens, is then:
@@
$\text{MMA}([X;X_{lr}])=\text{Softmax}\left(\frac{\left([Q(X),\tilde{Q}(X_{lr})]\right)([K(X),\tilde{K}(X_{lr})]^{T})\cdot M}{\sqrt{d}}\right)[V(X),\tilde{V}(X_{lr})],$
@@
$L_{\textrm{FM}}(\theta;C_{T},X_{lr})=\mathbb{E}_{t\sim\mathcal{U}(0,1),X_{t}\sim p(X|t,C_{T},X_{lr})}\left[\|f_{\theta}(X_{t},t,C_{T},X_{lr})-u_{t}(X_{t};C_{T},X_{lr})\|^{2}\right].$
@@
Fused-Kernel Adaptation. By default, X is reshaped into 1D token sequences by a raster-scan of the 2D latent grid for transformer inputs, which results in a sparse, non-contiguous attention mask that is suboptimal for GPU execution. We address this by re-permuting the tokens in $X$ and $X_{lr}$ in a Hilbert-Curve order as shown in Figure 2. This clusters all tokens belonging to a single local window into a dense, contiguous block in memory. We also construct attention mask $M$ for allowed interactions: text attends to itself and $X$ , $X$ attends locally and to both text and $X_{lr}$ regions, $X_{lr}$ attends to text and itself locally. To achieve acceleration, we follow FlashAttention’s Dao et al. (2022); Dao (2023) tiling strategy, which allow us to skip computation for entire blocks of the attention matrix masked out. Following Zhang et al. (2025c; e), we adapt the implementation to SageAttention Zhang et al. (2025d; c) yielding additional improvements in speed and VRAM efficiency. The pseudocode for our Block Sparse FlashAttention implementation is provided in the supplementary materials.

### v2

# UltraGen: Efficient Ultra-High-Resolution Image Generation with Hierarchical Local Attention

@@ Abstract
Ultra-high-resolution text-to-image generation is increasingly vital for applications requiring fine-grained textures and global structural fidelity, yet state-of-the-art text-to-image diffusion models such as FLUX and SD3 remain confined to sub 2MP (< $1K\times 2K$) resolutions due to the quadratic complexity of attention mechanisms and the scarcity of high-quality high-resolution training data. We present UltraGen, a novel framework that introduces hierarchical local attention with low-resolution global guidance, enabling efficient, scalable, and semantically coherent image synthesis at ultra-high resolutions. Specifically, high-resolution latents are divided into hardware aligned fixed-size local windows to reduce attention complexity from quadratic to near-linear, while a low-resolution latent equipped with scaled positional embeddings injects global semantics as an anchor. A lightweight LoRA adaptation bridges global and local pathways during denoising, ensuring consistency across structure and detail. To maximize inference efficiency and achieve scalable ultra-high-resolution generation, we repermute token sequence in window-first order, so that the GPU-friendly dense local blocks in attention calculation equals to the fixed-size local window in 2D regardless of resolution. Together UltraGen reliably scales the pretrained model to resolutions higher than $8K$ with more than $10\times$ speed up and significantly lower memory usage. Extensive experiments demonstrate that UltraGen achieves superior quality while maintaining computational efficiency, establishing a practical paradigm for advancing ultra-high-resolution image generation.
Figure 1: Ultra-high-resolution images generated by UltraGen with different base models (FLUX.1 and FLUX.2) at $8K\times 8K$, $4K\times 10K$, $8K\times 4K$, $6K\times 6K$, $4K\times 4K$, $4K\times 2K$, $2K\times 3K$. Zoomed-in regions highlight the fine-grained details preserved at these scales (See supplementary materials for original PDF). Notice that the pretrained FLUX 1 and FLUX 2 are only capable of generating sub 2MP (< $1K\times 2K$) and sub 4MP (< $2K\times 2K$) images.
@@ 1 Introduction
Ultra-high-resolution image generation is becoming increasingly crucial for both creative and practical applications. From digital art and advertising to scientific visualization and virtual environments, users demand outputs with fine-grained textures, faithful structures, and seamless details at $4K$ resolution or higher. However, despite the rapid progress of text-to-image (T2I) diffusion models [11, 40, 7, 35, 34, 54, 26, 6, 25, 2, 4, 50, 51], most state-of-the-art systems remain confined to resolutions below $1K\times 1K$. This limitation fundamentally restricts their usability in scenarios where ultra-high fidelity is paramount.
@@
The root of this challenge lies in two interdependent bottlenecks. First, scaling model capacity and datasets to higher resolutions requires tremendous resources, as high-quality 2K2K8K8K training data are scarce and expensive to curate. Second, attention-based diffusion architectures suffer from quadratic growth in token complexity with respect to image resolution, making naive scaling computationally prohibitive. For example, generating a $4K$ image would require tens of thousands of tokens, leading to impractical memory and runtime costs. As a result, existing approaches either resort to costly retraining on synthetic high-resolution datasets [19, 29, 38, 46, 69, 60], or adopt training-free upscaling methods [16, 36, 13, 30, 48, 44, 22, 21, 3] that often compromise efficiency and stability. Consequently, the field still lacks a solution that delivers both ultra-high-resolution fidelity and computational scalability.
@@
In this work, we introduce UltraGen, a novel framework that rethinks ultra-high-resolution synthesis through the lens of efficiency and scalability. Inspired by the way artists construct large-scale murals, we decompose the generation task hierarchically: local windows capture fine textures through lightweight attention, while a low-resolution guidance image preserves global semantic structure. Technically, UltraGen introduces three innovations. First, a hierarchical local attention mechanism restricts computation to fixed-size windows, reducing quadratic cost to near-linear scaling while yielding over $10GB$ memory savings and more than $10\times$ faster inference compared to dense attention. Second, a global guidance pathway employs a low-resolution latent with scaled RoPE positional anchors to maintain long-range dependencies and ensure semantic coherence across windows. Third, a parameter-efficient joint denoising framework integrates global and local pathways via LoRA-adapted projections trained only on $256\times 256$$1K\times 1K$ data, enabling direct scaling to $4K$ or higher without requiring any native high-resolution training. These design choices translate into two key outcomes: (1) scalable fidelity, achieved by hierarchically fusing global composition with local detail to extend beyond pretrained resolutions, and (2) computational efficiency, enabled by reducing computation, memory, and runtime while reusing pretrained weights with lightweight adaptation.
@@
Extensive experiments validate these contributions: UltraGen reliably scales to $4K\times 4K$ and higher resolution synthesis with commodity training resolutions, achieving sharper textures, richer details, and stronger global coherence than existing approaches. Quantitatively, it surpasses dense attention baselines with over $10\times$ faster inference and lower memory usage, and it matches or outperforms state-of-the-art methods trained with native 4K data across FID, IS, and CLIP Score. Together, these results highlight UltraGen as a practical and effective solution for advancing ultra-high-resolution text-to-image generation.
@@
  • We demonstrate substantial efficiency improvements ($10\times$ speedup in $8K$, less memory usage) while achieving state-of-the-art quality at $4K,8K$ and higher, offering a practical and generalizable solution for real-world ultra-high-resolution generation.
@@ 2 Related Work
Text-to-Image Generation. Recent years have witnessed rapid progress in text-to-image (T2I) generation [11, 40, 7, 35, 34, 54, 26, 6, 25, 2, 4, 50, 51, 14, 23], where diffusion models now achieve near-photorealistic synthesis at resolutions up to $1K\times 1K$. The dominant paradigm relies on cross-attention mechanisms and DiT architecture [34], as seen in the PixArt series [7, 6, 5], or multi-modal diffusion transformers (MMDiT) such as FLUX.1.0-dev [2]. These architectures have established strong foundations for controllable and high-quality T2I synthesis. However, despite their success, existing methods struggle to scale beyond pretrained resolutions due to quadratic attention costs and the lack of large-scale high-resolution data. Our work builds directly on this line of research, but departs from the prevailing direction by focusing on resolution scalability and computational efficiency rather than retraining larger models.
@@
High-resolution image synthesis. Real-world applications increasingly demand resolutions of $4K$ or higher, sparking intense research interest in pushing beyond the $1K\times 1K$ barrier. Several works, including PixArt-$\Sigma$ and SANA 1.5 [51], have achieved near-$4K$ synthesis through extensive high-resolution pretraining, while others [19, 29, 38, 52, 46, 69, 60, 57, 56] pursue fine-tuning or training-from-scratch approaches using curated datasets. For example, [60] used wavelet supervision to enhance detail clarity, and [57] proposed lightweight fine-tuning for adapting to higher resolutions. Although effective, these methods remain constrained by the scarcity of high-quality high-resolution data and substantial GPU requirements, and since the lack of high quality data as lower resolutions, the performance sometimes degrades and sensitive to aspect ratio [56, 60]. More recent training-free strategies [16, 36, 13, 30, 48, 44, 22, 21, 3, 67, 68] avoid data collection by leveraging pretrained models directly. For instance, I-Max [13] aligns high-resolution flow with low-resolution manifolds, and HiFlow [3] introduces low-resolution initialization to guide denoising. While promising, these approaches often inherit significant runtime and memory overhead, limiting accessibility for broader use. Other recent methods, such as HiDiffusion [67] and FreCaS [68], both target UNet-based architectures. HiDiffusion is training-free and regularizes the denoising trajectory but still struggles to preserve long-range consistency, while FreCaS employs a coarse-to-fine strategy to synthesize only high-frequency content, which reduces the long-range dependency burden but requires careful stage alignment and CFG tuning. In contrast, our method explicitly decouples fine-detail generation from global semantic reasoning using local-window attention and low-resolution guidance, achieving stable and efficient ultra-high-resolution synthesis without multi-stage repair or delicate parameter balancing.
Attention Acceleration. As image and video resolutions increase, the quadratic complexity of attention becomes the dominant bottleneck. Works including the SANA series [50, 51, 70] leverage linear attention to reduce complexity. While such methods achieve satisfactory performance, the non-injective property and loss of attention spikiness [17, 33, 65] inherent in linear attention lead to confusion and inconsistent local details in real-world scenarios. For softmax attention, system-level optimizations such as Flash Attention [9, 8, 43] exploit GPU features for faster execution, while quantization [62, 61] and sparsity-based designs [10, 27] reduce computational load. Architectural innovations, such as LongFormer [1] and SwinFormer [31], employ local attention patterns, and more recent works [24, 63, 53, 49, 55, 58, 59, 66, 28, 39] propose block sparsification or compression strategies for diffusion transformers. Although these techniques yield noticeable acceleration, the gains remain insufficient for ultra-high-resolution synthesis, and compression often risks degrading fine-grained fidelity. Methods like [28, 39, 66] aim to achieve faster generation while preserving local image quality using local neighbor attention. However, none of them fully maintain long-range dependencies: CLEAR [28] and GRAT [39] require increasingly large local windows/grouping as resolution increases, while STA [66] sacrifices some long-range consistency, also [28, 66] do not support ultra-high-resolution generation without other techniques like SDEdit [32]. In contrast, our method explicitly maintains long-range consistency at low resolution through hierarchical local-window attention, achieving stronger acceleration ($4\times$) and better detail preservation when scaling to 4K and higher, without relying on excessively large local windows or compromising global structure.
@@
Figure 2: Schematic of UltraGen’s attention block modifications. The left column illustrates that high-resolution image latents $X$ are partitioned into local windows (red) and flattened in window-first order that each window attends to itself. Simultaneously, a low-resolution guidance latent $X_{lr}$ (yellow) provides global context to each window via position scaling. The right column shows the joint-denoising process and the attention kernel, that $X,X_{lr}$ are processed together with LoRA applied on the $X_{lr}$ part. The attention mask (upper right) enforces local and guidance-specific interactions and tile skipping in attention calculation to enable efficient and coherent ultra-high-resolution generation. No additional high-resolution training data are needed.
@@ 3 Method
Figure 2 illustrates the overall framework of UltraGen, highlighting the key modifications introduced in the attention blocks to support ultra-high-resolution generation. High-resolution image latents $X$ are partitioned into local windows (red grids), where tokens attend only within their window to capture fine-grained detail efficiently. A low-resolution guidance latent $X_{lr}$ (yellow), enhanced with scaled positional anchors, injects global semantics and preserves long-range consistency across windows. The attention mask is designed to enforce both local and guidance-specific interactions while also enabling tile-skipping to avoid unnecessary computation. Together, these mechanisms allow UltraGen to generate coherent and detail-rich $4K$ images with near-linear complexity and without requiring any native high-resolution training data. This schematic provides the foundation for the following technical components.
The Multi-Modal Diffusion Transformer (MMDiT) Preliminaries. The MMDiT employed in state-of-the-art models [2, 4, 14, 23], represents the current benchmark architecture for text-to-image generation. MMDiT processes two distinct token modalities: a noisy image token sequence $X\in\mathbb{R}^{N\times d}$ and a text token sequence $C_{T}\in\mathbb{R}^{M\times d}$. MMDiT processes image and text tokens through a unified Multi-Modal Attention (MMA) mechanism. Specifically, image and text tokens are concatenated as $C_{T};X]$ and processed via self-attention.
Within the MMA framework, spatial information is encoded using Rotary Position Embedding (RoPE) [45], which applies rotational transformations to capture relative positional relationships. This is mathematically represented as: $X_{i,j}=X_{i,j}\cdot\text{Rot}(i,j)$, where $\text{Rot}(i,j)$ denotes the rotation matrix corresponding to position $(i,j)$ in the 2D spatial grid. The MMA mechanism $\text{MMA}([C_{T};X])$ is formally defined as:
$\text{Softmax}\left(\frac{\left([Q_{T}({C_{T}}),Q(X)]\right)\\ ([K_{T}(C_{T}),K(X)]^{T})}{\sqrt{d}}\right)[V_{T}(C_{T}),V(X)]$
@@
Efficient Local Window Attention with Window-First Permuted Token Sequence. Transformer-based generative models rely on self-attention, whose computational and memory costs scale quadratically with the number of tokens. For an image of size $H\times W$, the number of spatial tokens is $N=H\cdot W$, leading to $O(N^{2})$ complexity. Scaling from $1024\times 1024$ ($4096$ tokens [^1] A VAE downsampling factor of 16 yields $\frac{1024}{16}\times\frac{1024}{16}=4096$ tokens, similar scaling for other resolutions.) to $4096\times 4096$ ($65536$ tokens) increases the cost by a factor of $256$, which is prohibitive even with optimized kernels such as FlashAttention [9, 8, 43].
@@
To overcome this limitation, we partition the high-resolution latent $X\in\mathbb{R}^{H\times W}$ into non-overlapping windows $x_{i}\in\mathbb{R}^{l\times l}$}, with window size $l$ bounded by the pretrained resolution (e.g., 1024). Self-attention is computed independently within each window, reducing complexity from $O(N^{2})$ to $O\left(\lceil\tfrac{H}{l}\rceil\cdot\lceil\tfrac{W}{l}\rceil\cdot(l^{2})^{2}\right)\approx O(N\cdot l^{2})$. In practice, we set $l=16$ (corresponding to $256\times 256$ windows) to strike a balance between accuracy and efficiency: larger windows bring little additional benefit in quality, while smaller windows underutilize GPU kernels, which are optimized for tile sizes around $128\times 128$ [^2] With 16× VAE downsampling, 256×256 windows yield 16×16 tokens, creating attention matrices that efficiently utilize GPU kernels. Smaller window size produces tokens less than the granularity of 128.. An ablation study in Section 5 further analyzes this trade-off. With this design, scaling to $4K$ resolution ($N=65536$ tokens) reduces the attention complexity from $O(65536^{2})$ to $O(65536\cdot 16^{2})$, yielding a $256\times$ reduction. To maintain boundary consistency, we also allow limited cross-window attention along adjacent regions. Moreover, because the per-window computation cost remains constant, total runtime scales nearly linearly with image resolution, making the generation process both efficient and memory-friendly via integration with the tiling strategy in FlashAttention [9, 8].
A key aspect of our approach is the window-first token permutation: we reorder the flattened image token sequence in a window-first fashion (instead of the default raster-scan order) so that each block in the attention map corresponds directly to a $256\times 256$ spatial window in the original image. This ensures that every attention window operates within a fixed RoPE positional range of $256\times 256$ which is within the relative range which the model was pretrained regardless of the image resolution, thereby enabling ultra-high-resolution generation without extrapolating beyond learned positional encodings.
@@
Maintaining Global Semantics via Low-Resolution Guidance. Partitioning a high-resolution image into local windows risks creating discontinuities and losing global semantic coherence. The key insight is that global long-range dependencies primarily influence the overall structure and layout of the generated image, but have minimal effect on fine local details. We generate a low-resolution guidance image $X_{lr}\in\mathbb{R}^{h\times w}$ to provide global context. The corresponding position indices are denoted as (m,n){(m,n)}, where $m\in\{0,1,\dots,h-1\}$ and $n\in\{0,1,\dots,w-1\}$. We define the scaling ratio as $\rho=\frac{H}{h}$ (empirically set to 4 for an optimal balance of performance and efficiency). We then scale the low-resolution image’s position indices by this ratio, mapping them to $(\tilde{m},\tilde{n})=(\rho\cdot m,\rho\cdot n)$. This effectively projects $X_{lr}$ to the same spatial scale as the high-resolution image $X$. Each token $X_{lr}[\tilde{m},\tilde{n}$] acts as an anchor point, providing contextual information. The resulting attention structure is as follows: each high-resolution window attends to its own local and adjacent tokens together with the corresponding scaled region in $X_{lr}$, while $X_{lr}$ tokens attend globally among themselves. In concrete terms, we formulate the attention mask $M$ by allowing the following interactions: text tokens attend to themselves and to $X$; tokens in X attend locally and both text and $X_{lr}$; and tokens in $X_{lr}$ attend to each other. This design ensures semantic consistency across spatially distant regions. The framework naturally supports recursive generation: the high-resolution output can serve as low-resolution guidance for an even higher resolution, enabling arbitrarily large-scale synthesis with stable memory and computation.
@@
Parameter-Efficient Joint Denoising. To integrate the low-resolution guidance $X_{lr}$, we concatenate it with the standard text–image sequence to form $C_{T},X,X_{lr}]$. Since $X_{lr}$ is still an image modality, we can reuse the pretrained VAE and DiT blocks for its processing. However, scaling the positional indices of $X_{lr}$ alters the frequency characteristics of the RoPE embeddings. To adapt to this change, we fine-tune only the query, key, and value projections ($Q,K,V$) that process $X_{lr}$ using LoRA [20], yielding adapted matrices $\tilde{Q},\tilde{K},\tilde{V}$. Importantly, the original $Q,K,V$ are still used for the high-resolution windows $X$, ensuring that pretrained generative capabilities for local content are preserved. The unified attention mechanism $\text{MMA}([X;X_{lr}])$, simplified by omitting text tokens, is then:
@@
$\text{Softmax}\left(\frac{\left([Q(X),\tilde{Q}(X_{lr})]\right)([K(X),\tilde{K}(X_{lr})]^{T})\cdot M}{\sqrt{d}}\right)[V(X),\tilde{V}(X_{lr})],$
@@
$L_{\mathrm{FM}}(\theta;C_{T},X_{lr})$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,1),\\ X_{t}\sim p(X\mid t,C_{T},X_{lr})\end{subarray}}\Big[\big\|f_{\theta}(X_{t},t,C_{T},X_{lr})-u_{t}(X_{t};C_{T},X_{lr})\big\|^{2}\Big].$
@@
Inference Acceleration with Low-Res Guidance. Since the low-resolution guidance image already encodes the global structure of the target output—differing primarily in high-frequency detail—we leverage it to warm-start the high-resolution sampling process. Specifically, we first upsample the low-resolution reference to match the target resolution, apply sharpening, and add noise scaled to an intermediate timestep $t$, yielding $X_{t}^{\mathrm{ref}}=\mathrm{Interpolate}(X_{\mathrm{lr}})$. The high-resolution latent is then initialized as
$X_{\mathrm{hr}}^{t}=\alpha\,X_{T}+(1-\alpha)\,X_{t}^{\mathrm{ref}},$ (1)
where $\alpha=\frac{t}{T}$ controls the noise ratio and $X_{T}$ denotes pure Gaussian noise. By injecting the low-frequency components directly, this initialization allows us to bypass the early denoising stages, substantially accelerating inference.